OpenStudy (anonymous):

What is the solution of log small7 (4x + 5) = 2? A) 2 B) 6 C) 9 D) 11 E) 14

ganeshie8 (ganeshie8):

\[\huge \log_7(4x+5) = 2\]

ganeshie8 (ganeshie8):

is the equation like this... ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\[\huge {\log_ba =x\\=> a = b^x}\]

ganeshie8 (ganeshie8):

\[\huge4x+5 = 2^7\] \[\huge4x+5 = 128\] now its easy.... ?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

are you sure thats right because i'm getting 30.75 and thats not one of the answers

OpenStudy (anonymous):

It's wrong. He gave you the right theorem, but used it incorrectly. It's: \[4x+5=7^{2}\]

ganeshie8 (ganeshie8):

RHS should be 49 its not 128

ganeshie8 (ganeshie8):

@Wired thanks for catching :)